In 1895, the first U.S. Open Golf Championship was held. The winner’s prize money was $150. In 2016, the winner’s check was $1,800,000.

What was the percentage increase per year in the winner’s check over this period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)



If the winner’s prize increases at the same rate, what will it be in 2040? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.

1,800,000=150*(1+r/100)^121

(1,800,000/150)^(1/121)=(1+r/100)

(1+r/100)=1.0807

r=1.0807-1

=8.07%(Approx).

b.A=1,800,000*(1.0807)^24

=$11597551.4(Approx).